OpenStudy (anonymous):

what is the factor form of 3x-12x-12

OpenStudy (unklerhaukus):

factor out three first ,

OpenStudy (anonymous):

is it 3x2?

OpenStudy (anonymous):

3x - 12x - 12 = - 9x - 12 = 3(- 3x - 4) = -3(3x + 4)